OpenStudy (anonymous):

need help with Algebra 2 (will FAN & MEDAL)

OpenStudy (anonymous):

write an equation of a parabola with the given focus and directrix. focus: (-2,0);directrix: x=4

OpenStudy (anonymous):

no one likes these conic section question but they are not that hard if you know a couple things do you know what this looks like?

OpenStudy (anonymous):

ok lets go slow, but not that slow \(x=4\) is a vertical line right (not horizontal like say \(y=4\))

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this is where you say "yes i know that "

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

lol i do know that

OpenStudy (anonymous):

so lets graph the line \(x=4\) and also plot the point \((-2,0)\)

OpenStudy (anonymous):



OpenStudy (anonymous):

NOW do you know what it looks like? opens left, right, up or down?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

left?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(0,-2)?

OpenStudy (anonymous):

next we need to know the vertex, which is half way between the focus and the directrix what is half way between them

OpenStudy (anonymous):

look at the picture and visualize the point half way between the focus and the directrix your answer should be a coordinate, not a number also not \((0,-2)\)

OpenStudy (anonymous):



OpenStudy (anonymous):

1,0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so now we know it looks like \[y^2=-4p(x-1)\] and that since general form is \[(y-k)^2=4p(x-h)^2\]and you have \((h,k)=(1,0)\)

OpenStudy (anonymous):

the minus sign because it opens to the left

OpenStudy (anonymous):

also that is the reason the \(y\) term is squared, not the \(x\) term all you need it \(p\)

OpenStudy (anonymous):

\(p\) is the distance between the focus and the vertex, or the vertex and the directrix, which you pretty much know already

OpenStudy (anonymous):

no \(p=3\)

OpenStudy (anonymous):

so \[y^2=-12(x-1)\]

OpenStudy (anonymous):

i see i made a typo above there is not square on the x term

OpenStudy (anonymous):

\[(y-k)^2=4p(x-h)\]

OpenStudy (anonymous):

the distance between \((-2,0)\) and \((1,0)\)

OpenStudy (anonymous):

or from \((1,0)\) to \((4,0)\) same distance

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[y ^{2}=-12x+12 \]

OpenStudy (anonymous):

\[x=\frac{ -1 }{ 12 } y^2+1\]

OpenStudy (anonymous):

thank you